is this systen of equations:
-4x-2y=-4
3x-y=13
independent,dependent, or inconsistent

Answer by ankor@dixie-net.com(22740)

-4x - 2y = 4
3x - y = 13
Use the 2nd equation for substitution
-y = - 3x + 13
Mult by -1
y = (3x-13)
:
Replace y with (3x-13) in the 1st equatio
-4x - 2(3x-13) = 4
-4x - 6x + 26 = 4
-10x = 4 - 26
-10x = -22
x = 22/10
x = 2.2
:
Find y
y = 3(2.2) - 13
y = 6.6 - 13
y = -6.4
:
You can check this solutions in the original equations